OpenStudy (anonymous):

Which line shows the first error in the solution? (1) 6(x+2)=24 (2) 6x+12=24 (3) 6x=12 (4) x = 12 x 6 (5) x=72

OpenStudy (texaschic101):

line 4 it should be : 6(x + 2) = 24 6x + 12 = 24 6x = 12 x = 12/6 x = 2

OpenStudy (anonymous):

Thanks

OpenStudy (texaschic101):

:)

OpenStudy (anonymous):

What is the correct solution to the equation? 4(x-3)=16

OpenStudy (texaschic101):

4(x - 3) = 16 --- distribute through the parenthesis 4x - 12 = 16 --- add 12 to both sides 4x = 16 + 12 4x = 28 --- divide both sides by 4 x = 28/4 x = 7

OpenStudy (anonymous):

Which line shows the first error in the solution? Three more than a number divided by 4 is –12. What is the number? (1) x/4+3=-12 (2) x/4 +3 -3= -12-3 (3) x/4= -9 (4) x/4 x 4= -9 x 4 (5) x= -36 And thanks again for helping me